OpenStudy (anonymous):

how can you prove without a caluculator sin44degrees+sin16degrees=sin76degrees

OpenStudy (shubhamsrg):

now sin a+ sin b = 2sin((a+b)/2) cos((a-b)/2) thus LHS =2 sin(30) cos (14) = cos 14 =sin (90-14) =RHS hence proved!! omg!! ;)

OpenStudy (anonymous):

 now put the values and get the answer.